OpenStudy (anonymous):

Solve this equation for x. Round your answer to the nearest hundredth. 1= ln(x+8) Use the property z=a^logaZ

OpenStudy (anonymous):

applying that property, you know that: \(e^{\ln x} = x\) .. for your question, change the bases into e on both sides, so it will look like this: \(\sf \large e^1= e^{ln(x+8)}\) now simplify this using the property

OpenStudy (anonymous):

Im still confused...could you explain the process??

OpenStudy (anonymous):

if \(\ln(\text{whatever})=1\) then \(\text{whatever}=e\)

OpenStudy (anonymous):

How do i go about finding the answer? I dont understand

OpenStudy (anonymous):

Would it be 2??

OpenStudy (owlet):

What is e^1?

OpenStudy (anonymous):

I dont know what e^1 is

OpenStudy (owlet):

then using the rule, what is e^ln(x+8) simplify to? it is like z=a^log a z, a= e z= x+8 , therefore, e^ln(x+8) is equal to?

OpenStudy (owlet):

any number raised by one will be equal to the number itself., so e^1 =e

OpenStudy (anonymous):

I still dont understand how to find the answer.

OpenStudy (owlet):

simplify the equation first before you can find the value of x

OpenStudy (anonymous):

Ok

OpenStudy (owlet):